Need to find the volume of a right circular cylinder and round 2 places. r-5, h-29.
(5) (29)
(25) (29)
3.14x50x29=4553
I can't figure out where the decimal goes and how to round it 2 places.

Answer by cleomenius(959)

First, we find the area
Area = pi*r^2
5^2 = 25 * pi =78.54
Now, multiply this by the height.
78.54*29 = 2277.66
==================================================================
To determine where the decimal goes, determine what the place where the whole integers would be after the procedure.
If doing by hand, it is best to line up your figures in a column.
@|@|@|3|.|1|4
@|*|5|0|.| (this should be 25)
@|@|2|9|.|
----------
4|5|5|3|.|
=================================================================
One thing you did wrong was to use 50 instead of 25 in your calculations.
3.14x50x29=4553.
=================================================================
Cleomenius.